In a study using 8 samples, and in which the population variance is unknown, the distribution that should be used to calculate confidence intervals is
a. a standard normal distribution.

b. a t distribution with 7 degrees of freedom.

c. a t distribution with 8 degrees of freedom.

d. a t distribution with 9 degrees of freedom.

Answer by jim_thompson5910(35256)

Sample size = n = 8
Because n < 30 and "the population variance is unknown", we need to use the T distribution.

The degrees of freedom are df = n-1 = 8-1 = 7.

This all points to the answer being choice B.